A bank offers 10% interest rate on saving accounts, what uniform annual amount should be deposited each year in order to accumulate $1,000,000 at the time of the twentieth annual deposit?
Answer #1

ANSWER:

FV = $1,000,000

I = 10%

N = 20 YEARS

A = ?

FV = A(F/A,I,N)

1,000,000 = A(F/A,10%,20)

1,000,000 = A * 57.275

1,000,000 / 57.275 = A

A = 17,459.62

SO THE UNIFORM AMOUNT THAT SHOULD BE DEPOSITED EACH YEAR IS $17,459.62